# The Euler class of the normal bundle of a Seifert fibration and horizontal foliations

### v1

@@ Abstract.
For Seifert fibred manifolds with orientable base orbifolds, we establish a necessary and sufficient condition for the Euler class of the normal bundle of the Seifert fibration to vanish. When the base orbifold is hyperbolic, we also provide a second proof of this condition from the perspective of discrete faithful representations of Fuchsian groups.
As an application, we present infinitely many Seifert fibred rational homology spheres that admit co-oriented taut foliations but none with vanishing Euler class. In the context of the $L$-space conjecture, these provide examples of rational homology spheres that admit co-oriented taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups yet none of the left-orders arise directly from the universal circle actions associated to co-oriented taut foliations.
@@
Key words: Seifert fibred manifold, normal bundle, Euler class, horizontal foliation
@@ 1. Introduction
In this article, we study the triviality of the normal bundle $\nu_{M}$ to the Seifert fibres of a Seifert fibration on a $3$-manifold $M$ with orientable base orbifold. We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. As an oriented $2$-plane bundle, $\nu_{M}$ is trivial if and only if its Euler class $e(\nu_{M})\in H^{2}(M)$ vanishes.
@@
We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §2.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
Here is our main result.
@@
Remark 1.2.
It is often easy to identify an initial $m\in\mathbb{Z}$ that satisfies the congruence equations $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Any other integer $m^{\prime}$ satisfying these congruences can be written as $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Hence when $n>0$, $e(M)\neq 0$ and $\chi(\mathcal{B})\neq 0$, the condition $m\,e(M)=\chi(\mathcal{B})$ in Theorem 1.1 can be replaced by
$\frac{\chi(\mathcal{B})}{e(M)}\text{ is an integer congruent to }m\;(\operatorname{mod}l)$
@@
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an oriented $S^{1}$-bundle over the torus or one of the following holds:
@@
Note that if $M$ is an oriented $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 2.2).
@@
Our initial motivation for studying the Euler class of $\nu_{M}$ comes from the study of horizontal foliations on Seifert fibred manifolds. By definition, a horizontal foliation is transverse to the Seifert fibration and as such, their tangent bundles are isomorphic to $\nu_{M}$. Hence, their Euler classes coincide, at least up to sign.
It is predicted by the $L$-space conjecture [3, 13] that if a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving left-orderability of the fundamental group in the presence of a co-orientable taut foliation is to use Thurston’s universal circle action. See [7, 4] for the details. However, this approach requires the Euler class of the taut foliation to vanish. Although it was expected that examples of $3$-manifolds admitting co-orientable taut foliations but none with vanishing Euler class should exist, no published examples were previously known. In this article, we settle the matter by presenting such examples within the family of Seifert fibred rational homology $3$-spheres. Our arguments are based on the following application of Theorem 1.1.
Theorem 1.5.
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
The simplest examples of manifolds in which every co-oriented taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot. In the following statement, we assume that $p$ and $q$ are nonzero coprime integers.
Theorem 1.6.
The $3$-manifold obtained by $p/q$-surgery along the right-handed trefoil knot admits a co-oriented taut foliation if and only if $p/q<1$. It admits one with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
We will present more examples of rational homology Seifert fibred manifolds that admit co-oriented taut foliations, but none with vanishing Euler class, in §4.3 (Examples 4.4). In fact, our computations in §4 suggests that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
@@
In Section 2 we explain our notation for Seifert manifolds and prove Theorem 1.1, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.3 in this section. In Section 3 we present a different perspective on Theorem 1.1 and Corollary 1.3 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.7. Finally, in Section 4 we prove Theorem 1.5 and Theorem 1.6 and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §4.1 and is Euclidean in §4.2. The Euclidean case of Theorem 1.7 is dealt with at the end of §4.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §4.3. Finally, in §4.4, we illustrate how the conditions of Theorem 1.1 can be used to verify the known fact that $e(\nu_{M})=0$ when $|H_{1}(M)|=1,2$.

### v2

# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

@@ Abstract.
In this article, we construct infinitely many (small Seifert fibred, hyperbolic and toroidal) rational homology $3$-spheres that admit co-orientable taut foliations, but none with vanishing Euler class. In the context of the $L$-space conjecture, these examples provide rational homology $3$-spheres that admit co-orientable taut foliations (and hence are not $L$-spaces) and have left-orderable fundamental groups, yet none of the left orders arise directly from the universal circle actions associated to co-orientable taut foliations.
The hyperbolic and non-Seifert toroidal examples are obtained from Dehn surgeries on knots in the $3$-sphere and use Heegaard Floer homology to obstruct the existence of a co-orientable foliation with vanishing Euler class. For the Seifert fibred case, we establish necessary and sufficient conditions for the Euler class of the normal bundle of the Seifert fibration to vanish. Moreover, when the base orbifold is hyperbolic, we also provide a second proof of this condition from the viewpoint of discrete faithful representations of Fuchsian groups.
@@
Duncan McCoy was partially supported by NSERC grant RGPIN-2020-05491 and a Canada Research Chair
@@
Key words: taut foliation, Knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, Euler class, horizontal foliation.
@@ 1. Introduction
1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
It is predicted by the $L$-space conjecture [4, 18] that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
$\rho_{\mathcal{F}}:\pi_{1}(M)\to\mbox{Homeo}_{+}(S^{1}).$
See [10] for the details. If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [8, Theorem 1.1] that the fundamental group of $M$ is left-orderable. However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes. Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to the following question.
Question 1.1.
If a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, does it admit one whose Euler class is zero?  [^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign. So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
Although it has been expected that the answer to the question is no, we do not know of any examples verifying this. In this article, we show that such examples occur quite frequently.
Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p\geq 0$.
Theorem 1.2.
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $M$ does not admit a co-orientable taut foliation with Euler class zero.
Thus the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no. For instance, Roberts has shown that if $K$ is a non-trivial fibred knot in the $3$-sphere, then $K(p/q)$, the manifold obtained by $p/q$-surgery on $K$, admits a co-orientable taut foliation for all $p/q$, $p/q<1$, or p/q>1p/q>-1 depending on whether the fractional Dehn twist coefficient of $K$ is zero, positive, or negative respectively [26]. The following corollary states the case when the fractional Dehn twist coefficient is zero. Examples of such knots include all hyperbolic fibred alternating knots [6, 2].
Corollary 1.3.
Suppose that $K$ is a non-trivial fibred knot in the $3$-sphere whose fractional Dehn twist coefficient is zero. Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
A similar conclusion holds for many non-fibred knots. Indeed, the $L$-space conjecture predicts that every irreducible rational surgery on a non-fibred knot admits a co-orientable taut foliation and this is known to be true in many instances. For instance, Roberts and Delman proved this for non-torus alternating knots and for Montesinos knots which are not $L$-space knots.
Examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations. For instance, the figure eight knot is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [27].
Corollary 1.4.
Suppose that $K$ is the figure eight knot. Then for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
By way of contrast, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot without some hypotheses on the slope $p/q$. For instance, it is known that $p$-surgery on the figure eight knot admits co-orientable taut foliations of zero Euler class for each integer $p$ [11, 14]. More generally, we have
Proposition 2.3.
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the $3$-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
Remark 1.5.
The line of reasoning used in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot, meaning a knot with positive almost $L$-space surgeries [9], and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero. This shows that for knots of arbitrarily large genus, constraints must be placed on the denominators of surgery slopes for Theorem 1.2 to hold.
Corollary 1.6.
Suppose that $K$ is a knot of genus one in the $3$-sphere. If $p$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable. Similarly, if $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
All the examples in Corollary 1.4 are hyperbolic. The simplest examples of Seifert fibred manifolds in which every co-orientable taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
Theorem 1.7.
Suppose that $K$ is the right-handed trefoil knot. Then $K(p/q)$ admits a co-orientable taut foliation and has a left-orderable fundamental group if and only if $p/q<1$. It admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
@@ 1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
In general, we denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §3.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
Theorem 1.8.
  1. (1)
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
We can also produce toroidal examples which are not Seifert fibred. Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$. Any finite rooted tree can arise as the JSJ graph of a knot. We say that a JSJ graph is a rooted interval when it is homeomorphic to an interval with root at an endpoint. Since non-integral surgeries on knots in $S^{3}$ are irreducible [12], the following corollary follows immediately from Theorem 1.2 and Theorems 3.13 and 4.3 of [3].
Corollary 1.9.
Suppose that $K$ is a fibred satellite knot whose $JSJ$ graph is not a rooted interval. Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ is toroidal, has a left-orderable fundamental group, and admits a co-orientable taut foliation, but none with zero Euler class.
Remark 1.10.
Fibred composite knots satisfy the corollary’s hypothesis. Taking $K$ to be a connected sum of torus knots yields examples which are graph manifolds.
1.2. Euler class of the normal bundle of a Seifert fibration
While Theorem 1.2 is proved using the knot Floer complex, the key fact in analyzing the Euler classes of co-oriented taut foliations on a Seifert fibred rational homology $3$-sphere $M$ is that, up to isotopy, all such foliations are horizontal; that is, transverse to the Seifert fibres of $M$ (Lemma 5.1). Hence, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$, while the latter makes sense even without the existence of a horizontal foliation.
We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. In §3 and §4, we determine the following necessary and sufficient conditions for its Euler class to vanish.
@@ 1.2. Euler class of the normal bundle of a Seifert fibration
Theorem 1.8 follows from Theorem 1.11 and Lemma 5.1.
@@
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an orientable $S^{1}$-bundle over the torus or one of the following holds:
@@
Note that if $M$ is an orientable $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 3.2).
@@
We will present more examples of rational homology Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3 (Examples 5.3). In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
@@
We prove Theorem 1.2 and Proposition 2.3 in §2. In §3 we explain our notation for Seifert manifolds and prove Theorem 1.11, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.12 in this section. In §4 we present a different perspective on Theorem 1.11 and Corollary 1.12 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.14. Finally, in §5 we prove Theorems 1.7 and 1.8, and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2. The Euclidean case of Theorem 1.14 is dealt with at the end of §5.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §5.3.
@@ Acknowledgements
2. Examples from surgeries on knots in the $3$-sphere
Theorem 1.2 is an immediate consequence of the following two lemmas. The first follows from Ozsváth and Szabó’s proof that a manifold with a co-orientable taut foliation is not an $L$-space [22].
Lemma 2.1.
Let $M$ be a rational homology $3$-sphere admitting a co-orientable taut foliation $\mathcal{F}$. Then $M$ has a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
Proof.
Let $M$ be a rational homology $3$-sphere that admits a co-orientable taut foliation $\mathcal{F}$. The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [22]. ∎
The second is an application of the mapping cone formula for calculating the Heegaard Floer homology of surgeries on a knot in the $3$-sphere [23, 24]. In conjunction with Lemma 2.1, this implies Theorem 1.2 since for a 3-manifold $M$ for which $H_{1}(M)$ has odd order, the unique spinc-structure satisfying $c_{1}(\mathfrak{s})=0$ is also the unique self-conjugate spinc-structure on $M$.
Lemma 2.2.
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere, where $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1. Then the unique self-conjugate $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
Proof.
Without loss of generality we may assume that $p/q>0$. Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex for $K$. Since |p/q|>2g(K)1|p/q|>2g(K)-1, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
$\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that [21, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K),$
where $\bm{A}_{{\rm red},\,k}(K)$ denotes the reduced part of
$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
However, $\bm{A}_{{\rm red},\,k}(K)=0$ for |k|>g(K)1|k|>g(K)-1 [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&0\leq i<qg(K)-q,\\ \bm{A}_{{\rm red},\,\lfloor\frac{i-p}{q}\rfloor}(K),&p-qg(K)<i<p,\\ 0,&qg(K)-q\leq i\leq p-qg(K).\end{cases}$
In particular, for each $|j|\leq g(K)-1$, there are precisely $q$ $\mathrm{spin}^{c}$-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})$ is isomorphic to $\bm{A}_{{\rm red},\,j}(K)$, and for all other $\mathrm{spin}^{c}$-structures, $HF_{\mathrm{red}}(M,\mathfrak{s})=0$. Thus the set
$S=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0\}$
has cardinality divisible by $q$. In particular, $|S|$ is even. Since conjugation of $\mathrm{spin}^{c}$-structures induces an isomorphism of Heegaard Floer homology, conjugation acts as an involution on $S$. This implies that there are an even number of self-conjugate $\mathrm{spin}^{c}$-structures in $S$. Since $p$ is odd, there is precisely one self-conjugate spinc-structure on $M$, so it follows that $S$ does not contain the self-conjugate spinc-structure on MM. [^2] An alternative line of argument is to use the fact that conjugation of spinc-structures corresponds to the involution $i\mapsto p+q-1-i$ (mod $p$) on $\mathbb{Z}/p$. For $q$ even, this implies that the self-conjugate spinc-structure is labelled by $s=\frac{p+q-1}{2}$. However since |p/q|>2g(K)1|p/q|>2g(K)-1, the above calculation shows that $HF_{\mathrm{red}}(M,\phi(s))=0$.
In the opposite direction, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot $K$ without some hypotheses on the slope $p/q$.
Proposition 2.3.
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the 3-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
Proof.
Let $K$ be a genus one knot and let $M$ be obtained by $p$-surgery on $K$. Without loss of generality we may assume that $p>0$. The proposition is deduced from Lemma 2.1 by showing that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and then verifying that $c_{1}(\mathfrak{s}_{0})=0$.
First we compute $HF_{\mathrm{red}}(M)$. For integer surgeries of slope $p\geq 2g(K)-1$, the mapping cone formula gives a labelling
(2.0.1) $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that [21, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,i+kp}(K).$
However, as $g(K)=1$, we have $\bm{A}_{{\rm red},\,k}(K)=0$ for $|k|>0$ [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,0}(K),&i=0,\\ 0,&0<i\leq p-1,\end{cases}$
giving that that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and that this is the spinc-structure corresponding to 0 under (2.0.1). We verify that $c_{1}(\mathfrak{s}_{0})=0$ by unpacking the labelling in (2.0.1). As in [23, Section 2.4] the map $\phi$ is obtained as follows. Take $X$ to be the trace of $p$-surgery on $K$. Let $\Sigma]\in H_{2}(X)$ denote that the homology class obtained by capping off the core of the 2-handle with a Seifert surface for $K$. If $\mathfrak{s}\in\mathrm{Spin}^{c}(M)$ is the restriction of $\mathfrak{t}\in\mathrm{Spin}^{c}(X)$, then $\phi(i)=\mathfrak{s}$, where $i$ satisfies
$\langle c_{1}(\mathfrak{t}),[\Sigma]\rangle+p\equiv 2i\bmod 2p.$
Thus the spinc-structure with $\mathfrak{s}_{0}=\phi(0)$ is the restriction of any spinc-structure $\mathfrak{t}_{0}$ on $X$ such that $\langle c_{1}(\mathfrak{t}_{0}),[\Sigma]\rangle\equiv p\bmod 2p$. Since the restriction map $H^{2}(X)\rightarrow H^{2}(M)$ is reduction mod $p$, this implies that $c_{1}(\mathfrak{s}_{0})=0$, as required. ∎
Remark 2.4.
As mentioned in the introduction, the line of reasoning in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.